Can you tell me a good way to solve linear equations.I'm really stuck.....I need an explaination that you would give someone in 2nd grade.My mom and my teacher have explained it over and over but I just don't get it!!!
Answer by Nate(3500)

The basic equation for a linear line is y=mx+b
if you get the equation in this form, you will know all about solving the line
while m is the slope, b is the y-intercept (the point that crosses the y-axis)
lets say you have the equation y=3x-3
your y-intercept is -3
your slope is 3/1
lets say you have the equation x+2y=4
you need to make is look like the formula
2y=-x+4
y=(-1/2)x+2
your y-intercept is 2
your slope is -1/2